OpenStudy (anonymous):

the label on a can of paint states that it will cover 64 square feet what expression will show the length of a square wall that one can of paint will cover

terenzreignz (terenzreignz):

This is a square, so the length of the sides must be equal. The area is 64 square feet. A = s^2 64 = s^2 s = 8 Not sure if this is what you meant, feel free to correct me :)

OpenStudy (anonymous):

the anwser choices are 64/2, 64^2, 64 (2), and 64 i cant do the symbol with it but it looks like a check almost

terenzreignz (terenzreignz):

oh then, the 64 with a check is the correct answer :) The sign that looks like a check is the square root (radical sign), I also don't know how to put it here (I'm slow) But the 64 with a check is the same as 8, because 8*8 = 64 :) Hope that helps

OpenStudy (anonymous):

ok i thought it was square root lol thanks alot

terenzreignz (terenzreignz):

No problem :)